Can you give me step by step how to solve Linear Equations and why you did what you did? Please do equations like -4(3r-2)=-4 and 3(h-9)+2h=3
Answer by stanbon(75887)

-4(3r-2)=-4
----
Apply the distributive law to the left side to get:
----
-12r+8 = -4
----
Subtract 8 from both sides to get:
----
-12r = -12
Divide both sides by -12 to get:
----
r = 1
====================
3(h-9)+2h=3
----
Apply the distributive law to the 1st term of the left side to get:
3h-27 + 2h = 3
-----
Combine "like" terms to get:
5h - 27 = 3
------
Add 27 to both sides to get:
5h = 32
------
Divide both sides by 5 to get:
h = 6.4
============
Cheers,
Stan H.
==============